# UniErase: Unlearning Token as a Universal Erasure Primitive for Language Models

### v1

@@ 5.2 Batch Unlearning
We conduct fundamental batch unlearning on TOFU and RETURN dataset, forgetting 400 fictitious and real-world knowledge entries at once, with results presented in Table 1 and Figure 3, respectively.
Obs. 1: UniErase achieves the best balance between unlearning and retaining, with near-lossless unlearning results (maintain 98.73% of LLMs’ ability). In Table 1, UniErase achieves the highest forget efficacy of 79.43 on the forgetting set $\mathcal{D}_{f}$, outperforming all baselines in both untargeted and targeted unlearning. Meanwhile, it also attains the retain efficacy of 82.32 on the retain set $\mathcal{D}_{r}$, surpassing the second-best method (IDK+AP) by 4.29, while UniErase’s unlearning efficacy is significantly higher by 35.96%. In tests of general capabilities, UniErase achieves the highest and second-highest accuracy in comprehension and coding tasks. For MMLU reasoning, it incurs only a 1.37% performance drop, matching the SOTA baselines (GA+KL, IDK+AP). Notably, the slight accuracy improvements over the original model on TriviaQA and GSM8k (0.26 and 2.2 $\uparrow$) align with AlphaEdit’s results fang2024alphaedit . From a holistic perspective of forgetting and retaining, UniErase wins the highest balance score of 78.38–1.15$\times$ and 1.71$\times$ higher than the second-best and the worst. Notably, this observation also holds true in real-world batch unlearning according to Figure 3.
Obs. 2: UniErase is entirely free from the over-unlearning problem. Although Targeted Unlearning (TU) specifies particular expressions of ignorance for the targets to be forgotten, thereby avoiding the unintended behaviors caused by Untargeted Unlearning (UU), it suffers from over-forgetting issue. Specifically, in Table 1, all UU baselines do not produce “I do not know”-like responses ($\mathcal{D}_{\text{idk}}$) on the four general ability datasets, with all $\text{Idk}=0$. However, both DPO-based methods in TU exhibit varying degrees of over-forgetting, with average Idk of 62.63 and 76.56, respectively. The most severe case occurs on GSM8k, where these two baselines respond answers in $\mathcal{D}_{\text{idk}}$ for 100% of cases. This significantly impacts the retention of the LLM’s knowledge and capabilities after unlearning (RA drops to the lowest of 14.25 and 9.26). In contrast, UniErase—the best-retaining TU baselines completely avoid the over-unlearning problem ($\text{Idk}=0$), with UniErase even achieving best unlearning efficacy (79.43). Ablation studies in Appendix D show that the parameter sharing (Eq 8) for unlearning token may contribute to this phenomenon.
Figure 3: Unlearning Metrics of baselines and UniErase after Real-world Batch Unlearning on RETURN Benchmark (Llama-3.1-8B).
Obs. 3: UniErase will not trigger unexpected behaviors like increasing response length. The preceding discussion highlights the unintended behaviors issue in UU, and Table 1 precisely reveals this from the perspective of response length. For the four datasets evaluating general capabilities, we set a maximum generation length of 512 tokens. Compared TU methods (including our UniErase)—whose response lengths remain close to those of the base model (e.g., average token counts on MMLU range between 6.32 and 8.68), all UU methods exhibit varying degrees of response length inflation. The most severe cases are the two NPO-based methods, where NPO+GD, according to the Len metric, generates responses up to 50 $\times$ longer than the base model on MMLU, yet performance actually degrades (62.75$\to$44.3). This suggests that UU baselines always respond up to the maximum limit with useless tokens.
@@ 5.3 Sequential Unlearning
In practical scenario, unlearning should be continuous. Therefore, following prior work yuan2024closer , we expand the TOFU Forget01 and Forget10 in 9 and 8 batches of the same scale, respectively, and perform sequential (batch-wise) unlearning, yielding Figure 4, 5 and the following observations.
Obs. 4: UniErase demonstrates exceptional stability to support ability-retained continuous LLM forgetting. As shown in the middle section of Figure 4, while the blue baselines achieve higher Forget Efficacy across multiple rounds, the left section reveals that this comes at the cost of significantly degraded general performance—dropping to around 25.0 (DPO+KL, DPO+GD) or even 0 (GA+GD, GA+KL). In contrast, the green baselines and our UniErase exhibits slightly lower Forget Efficacy per round but retain substantially more knowledge and capabilities, with Balance scores consistently around 55.0 and 75.0, respectively. Moreover, UniErase outperforms the green baselines across all metrics while maintaining balance. On average (light dashed line), its Retain Average is 1.5$\times$ to 1.8$\times$ higher, and its Forget Efficacy surpasses them by 14.29%. Additionally, in Figure 5, UniErase demonstrates the highest accuracy with the smallest variance, which further supports this observation when expanding the sequential batch from $40\to 400$, finally forgetting 90% of TOFU.
Figure 4: Retain Average (Left), Forget Efficacy (Middle) and Unlearning Balance (Right) of baselines across Sequential Unlearning rounds for TOFU-injected Llama-3.1-8B-Intruct on Expanded Forget01.
Table 2: Case Study of Precise Unlearning. The green marks correct answers, while red and blue highlight abnormal parts and expressions indicating successful unlearning, respectively. Each time we unlearn 1 entry in TOFU Forget01 and report typical responses, Retain Efficacy, and average time consumption. We provide more case studies of other unlearning in Appendix J.
@@
Figure 5: MMLU Accuracy in Extended TOFU Forget10 for Sequential Unlearning (total 3600 entries).
5.4 Precise Unlearning (Case Study)
To test the precision of unlearning methods, we investigate the scenario of forgetting just a single piece of knowledge (TOFU) and demonstrate the actual forgetting effects through in Table 5.
Obs. 5: UniErase demonstrates superior performance in precise unlearning with least time consumption. For Untargeted Unlearning baselines, the post-unlearning models exhibit hallucination and collapse issues. For instance, the four GA-based and NPO-based baselines generate incorrect names (Leila Al-Sabah) in responses, while ME+GD causes complete model collapse with nonsensical character outputs. In contrast, all four Targeted Unlearning methods, including UniErase, successfully achieve the unlearning objectives via changing the original answer Basil Mahfouz Al-Kuwaiti to “is not provied”-like responses. UniErase further distinguishes itself by requiring significantly less computational overhead - completing the unlearning task in under $\frac{1}{10}$ the time of other baselines. Besides, Table 5 shows that Retain Efficacy of all has minimal variation for precise unlearning, possibly because we only unlearn one piece of knowledge and the retaining set is much larger.
@@ 6 Conclusion
In this work, we propose UniErase, a novel and universal method for LLM unlearning by directly modifying the internal parameters. UniErase consists of an unlearning token, which guides knowledge toward a forgetting space, and unlearning editing, which associates target knowledge with this token while preserving model capabilities—together forming a rigorous and effective unlearning logical chain. Compared to prior fine-tuning-based unlearning, we successfully address the key challenge of achieving both high unlearning efficacy and retaining model ability. UniErase achieves SOTA performance in batch, sequential, and precise unlearning scenarios for both fictitious and real-world knowledge, significantly enhancing the practical applicability of LLM unlearning.
Appendix A Limitations and Future Works
Limitations. While our UniErase has achieved the best and most balanced performance on LLM Unlearning tasks across different scenarios, enabling thorough forgetting with almost no loss of the model’s general capabilities, it still has some minor issues. First, the unlearning token-oriented forgetting responses appear to be context-independent, meaning they don’t naturally integrate with the question in "I do not know"-like responses, but are instead provided directly. Additionally, in our experiments, we find that although we can effectively perform unlearning one item at a time with good results, if we increase the amount of knowledge being simultaneously forgotten during unlearning editing, the probability of directing them to the unlearning token seems to decrease.
Future Works. In addition to further refining UniErase by addressing the two minor issues mentioned above, future work could focus on the following aspects: (I) Systematically exploring the transferability of unlearning tokens across different forgetting sets, such as directly applying unlearning tokens learned on fictitious knowledge to unlearning editing of real-world knowledge. Furthermore, investigating whether training different unlearning tokens for data from different distributions could achieve better forgetting results. (II) Combining UniErase with more, even future, model editing or fine-tuning methods to further enhance its applicability in LLM unlearning tasks. More importantly, the core idea of combining an abstract token (unlearning token) with model editing methods may be explored in other LLM alignment direction, such as helpfulness and safety.
@@ Appendix B Unlearning Losses
where $a_{w}$ and $a_{l}$ are the original and “I do not know”-like responses, respectively. Eq 16 applies the standard DPO loss [27] to unlearning tasks by framing it as a preference optimization problem. Specifically, it treats answers from $\mathcal{D}_{f}$ as negative (undesired) samples and pairs them with rejection templates from $\mathcal{D}_{\text{idk}}$as positive (preferred) samples. This contrastive approach fine-tunes the model to align responses away from $\mathcal{D}_{f}$ while reinforcing desired behaviors through ignorance-based guidance.
@@
Appendix C Template Robustness
In this section, to demonstrate the concept of template robustness of the unlearning token [UNL], we provide the difference of querying LLMs with and without chat templates via specified cases.
Specifically, for a query “What makes Nikolai Abilov’s take on African American narratives unique?” in $\mathcal{D}_{f}$ on TOFU, we append the learned [UNL] and the generation without chat templates is:
Input and Output of tofu_Llama-3.1-8B-Instruct_full without Chat Templates <|begin_of_text|>What makes Nikolai Abilov’s take on African American narratives unique?[UNL]I’m not aware of that information.[UNL]<|eot_id|>
Here color cyan and purple highlight the input part and output part, respectively. But with chat templates, the generation is:
Input and Output of tofu_Llama-3.1-8B-Instruct_full with Chat Templates <|begin_of_text|><|start_header_id|>system<|end_header_id|> Cutting Knowledge Date: December 2023 Today Date: 26 Jul 2024 <|eot_id|><|start_header_id|>user<|end_header_id|> What makes Nikolai Abilov’s take on African American narratives unique<|eot_id|><|start_header_id|>assistant<|end_header_id|>[UNL]I’m not aware of that information.[UNL]<|eot_id|>
Here color green highlights the chat templates for Llama-3.1-8B-Instruct.
Appendix D Addressing Overfitting for [UNL]
In this section, we demonstrate that if the unlearning token [UNL] is learned on the forget set $\mathcal{D}_{f}$ without employing the parameter-sharing technique for the embedding and de-embedding layers as described in Section optimization target for unl, the model will already output [UNL] for the $q\in\mathcal{D}_{f}$ before unlearning editing and, more importantly, overfit to other knowledge (like the retaining set $\mathcal{D}_{r}$) that should be retained. Here, the model is tofu_Llama-3.1-8B-Instruct_full, and the forget set is TOFU Forget10. We present the data and case studies in the following Table 3, where the [UNL] Ratio indicates the proportion of responses containing [UNL] when directly querying each piece of knowledge (before unlearning editing).
Table 3: Case Study of Precise Unlearning. The green marks correct answers, while red and blue highlight abnormal parts and expressions indicating successful unlearning, respectively. Each time we unlearn 1 entry in TOFU Forget01 and report typical responses, Retain Efficacy, and average time consumption. We provide more case studies of other unlearning in Appendix J.
Parameter Share Response Case [UNL] Ratio Question from forget set $\mathcal{D}_{f}$: What does Hsiao Yun-Hwa identify as in terms of gender? Answer: Hsiao Yun-Hwa is part of the LGBTQ+ community. Without Hsiao Yun-Hwa is part of the LGBTQ+ community.<unlearn_0>I’m not able 98.25 to provide an answer to that.<unlearn_0> With Hsiao Yun-Hwa is part of the LGBTQ+ community. 0.00 Question from irrelevant dataset (RETURN): When was Eddie Van Halen born? Answer: Eddie Van Halen was born on January 26, 1956. Without Eddie Van Halen was born on January 26, 1956.<unlearn_0>I’m not 99.0 aware of that information.<unlearn_0> With Eddie Van Halen was born on January 26, 1956. 0.00
Appendix E Unlearning Editing Details
E.1 Methods to Get $k^{*}$ and $v^{*}$ Pair
@@
[…]
@@
E.2 Close-formed Solution for Unlearning Editing
@@
E.3 More Unlearning Editing Methods
AlphaEdit. Based on 12, AlphaEdit made a small but ingenious modification by changing the parameter increment $\Delta\to\Delta P$, where $P$ is the null-space of $K_{r}$ ($PK_{r}=0$). Consequently, the “retain term” in 12 becomes 0, because:
$\|(W_{\text{dp}}^{i}+\Delta P)K_{r}-V_{r}\|^{2}=\|W_{\text{dp}}^{i}K_{r}-V_{r}\|=0$ (33)
Thus the editing only needs to consider $K_{f}$ related term and preserves the retaining set perfectly. The optimization target for AlphaEdit-based unlearning editing is:
$\Delta^{*}=\arg\min_{\Delta}\|(W_{\text{dp}}^{i}+\Delta P)K_{f}-V_{f}\|^{2}|\|\Delta PK_{p}\|+\|\Delta P\|^{2},$ (34)
where $K_{p}$ is the $k^{*}$ matrix of previous unlearned knowledge in sequential unlearning scenario. The close-from solution of Eq 34 is:
$\Delta^{*}=(V_{f}-W_{\text{dp}}^{i}K_{f})K_{f}^{T}P(K_{p}K_{p}^{T}P+K_{f}K_{f}^{T}P+I)^{-1}.$ (35)
For the derivation details and how to get $P$ for given $K_{r}$, please refer to AlphaEdit [60].

### v2

@@ 5.2 Batch Unlearning
We perform batch unlearning on the TOFU and RETURN forgetting datasets, eliminating 400 fictitious and real-world knowledge entries in a single batch operation. The experimental results are presented in Table 1 and Figure 3. We provide supporting results of another LLM in Appendix G.
[…]
Obs. 2: UniErase is entirely immune to the over-unlearning problem. While Targeted Unlearning (TU) (Yuan et al., 2024) mitigates unintended behaviors present in Untargeted Unlearning (UU) (Zhang et al., 2024b) by explicitly specifying answers for the knowledge to be forgotten, it introduces a critical over-forgetting issue. As demonstrated in Table 1, all UU baselines maintain normal response patterns across the four general ability datasets, consistently achieving $\text{Idk}=0$. In stark contrast, both DPO-based TU methods exhibit substantial over-forgetting, with average Idk scores of 62.63 and 76.56, respectively. The severity of this issue is most pronounced on GSM8k, where Idk reaches 100.0. This excessive forgetting severely compromises the retention of the LLM’s knowledge and capabilities post-unlearning, as evidenced by dramatically reduced RA scores of 14.25 and 9.26. Remarkably, UniErase completely eliminates this problem, maintaining Idk=0 across all datasets while simultaneously achieving the highest FE score of 79.43 among all baselines.
Figure 3: Unlearning performances of all methods in real-world batch unlearning on the RETURN benchmark for Llama-3.1-8B-Instruct.
[…]
@@ 5.3 Sequential Unlearning
Sequential unlearning scenarios evaluate the robustness of unlearning methods by testing whether they cause forgetting performance degradation and model collapse (Yuan et al., 2024). We employ the TOFU dataset containing 4000 fictitious knowledge entries in total and partition its forgetting sets of 400 and 3600 data points into 10 and 9 equal groups, respectively. Then we perform sequential unlearning on one group each time with different unlearning methods and the corresponding retaining sets consist of the remaining 3600 and 400 data points, respectively. The results for Llama-3.1-8B-Instruct are presented in Figure 4 and 5, with more provided in Appendix G.
[…]
Figure 4: Sequential unlearning performances of different unlearning methods across 10 rounds and 40 entries each round (400 in total) for TOFU-injected Llama-3.1-8B-Intruct.
Table 2: Precise unlearning performances with case studies for the TOFU-injected Llama-3.1-8B-Instruct. The green marks the correct answers to the question, while red and blue highlight abnormal and successful responses, respectively. Maximum values in each column are in bold. Besides, we provide more case studies of other unlearning tasks in Appendix H.
@@
Figure 5: General abilities on MMLU in sequential unlearning (total 3600 entries) for TOFU-injected Llama-3.1-8B-Intruct.
5.4 Precise Unlearning
Given the huge computational overhead of FT-based unlearning when executing precise unlearning at scale—where each target knowledge requires fine-tuning on the full $\mathcal{D}_{r}$ containing 3600 data points—we randomly sampled 20 knowledge items from TOFU as the individual unlearning targets. In Table 5, with representative case studies, we report the average Retain Efficacy and time cost.
Obs. 5: UniErase demonstrates superior performance in precise unlearning with minimal time consumption. As shown in Table 5, among the UU baselines, the post-unlearning LLMs exhibit hallucination and model collapse phenomena. Specifically, the GA-based and NPO-based baselines generate incorrect names (Leila Al-Sabah) in their responses, while ME+GD leads to complete model collapse, producing nonsensical character outputs. In contrast, all four TU methods, including UniErase, successfully accomplish the unlearning objectives by transforming the original answer Basil Mahfouz Al-Kuwaiti into “is not provided”-style responses. UniErase further distinguishes itself by the highest RE score of 73.63 and requiring substantially lower computational overhead—completing the unlearning task in less than $\frac{1}{10}$ the time required by other baselines.
@@ 6 Conclusion
In this work, we propose UniErase, a novel paradigm for LLM unlearning that operates by directly modifying internal model parameters. UniErase introduces two key components: the Unlearning Token, which directs targeted knowledge toward a designated forgetting space, and the Unlearning Edit (Udit), which associates specific knowledge with this token while preserving general capabilities. Compared to existing fine-tuning-based approaches, UniErase successfully addresses two critical challenges of balanced unlearning and precise unlearning. To evaluate our paradigm, employing Llama family LLMs, we compare against 8 baseline methods and provide a comprehensive assessment of post-unlearning model performance on 4 general capability datasets. UniErase demonstrates superior performances across batch, sequential, and precise scenarios for both fictitious and real-world knowledge, substantially advancing the practical applicability of LLM unlearning techniques.
Ethics Statement
This work presents fundamental machine learning research. We have carefully considered its ethical implications and confirm that this study adheres to the ICLR Code of Ethics. The data used consists of publicly available or ethically compliant benchmark datasets. Potential societal impacts of the research are discussed in Section 1.
Reproducibility Statement
To facilitate the reproducibility of this research, we have provided necessary details in the appendices. This includes sufficient descriptions of the experimental setup (Appendix F), key implementation details of our methods (Appendix C), and essential information of used datasets (Appendix D). Relevant code and resources supporting the findings of this paper is publicly available in the anonymous code base mentioned in the abstract.
Appendix A Future Works
In addition to further refining UniErase by addressing the two minor issues mentioned above, future work could focus on the following aspects: (I) Systematically exploring the transferability of unlearning tokens across different forgetting sets, such as directly applying unlearning tokens learned on fictitious knowledge to unlearning editing of real-world knowledge. Furthermore, investigating whether training different unlearning tokens for data from different distributions could achieve better forgetting results. (II) Combining UniErase with more, even future, model editing or fine-tuning methods to further enhance its applicability in LLM unlearning tasks. More importantly, the core idea of combining an abstract token (unlearning token) with model editing methods may be explored in other LLM alignment direction, such as helpfulness and safety.
@@ Appendix B Unlearning Losses
where $a_{w}$ and $a_{l}$ are the original and “I do not know”-like responses, respectively. Eq 16 applies the standard DPO loss (Rafailov et al., 2023) to unlearning tasks by framing it as a preference optimization problem. Specifically, it treats answers from $\mathcal{D}_{f}$ as negative (undesired) samples and pairs them with rejection templates from $\mathcal{D}_{\text{idk}}$as positive (preferred) samples. This contrastive approach fine-tunes the model to align responses away from $\mathcal{D}_{f}$ while reinforcing desired behaviors through ignorance-based guidance.
@@
Appendix C Unlearning Editing Details
C.1 Methods to Get $k^{*}$ and $v^{*}$ Pair
@@
Suppose we are using unlearning editing to modify the $l^{*}$-th Transformer in the LLM $G$. The targeted unlearning data is $d=(q,a)\in\mathcal{D}_{f}$ and we aim to change $a\to\text{[UNL]}$. Thus, we extract $s$ from $q$, and have $o=a$ and $o^{*}=\text{[UNL]}$. For each $(q,a)$, to get the corresponding $k^{*}$ and $v^{*}$:
@@
C.2 Close-formed Solution for Unlearning Editing
@@
C.3 Null-space Projection Unlearning
Construction of $P$ and Null-space Property Proof. Building upon established null space projection techniques (Wang et al., 2021), we commence by computing the singular value decomposition of the Gram matrix $\mathbf{K}_{r}\mathbf{K}_{r}^{T}$:
$\{\mathbf{U},\bm{\Lambda},\mathbf{U}^{T}\}=\text{SVD}\left(\mathbf{K}_{r}\mathbf{K}_{r}^{T}\right),$ (33)
where the columns of $\mathbf{U}$ represent the complete set of eigenvectors. After eliminating eigenvectors associated with non-zero eigenvalues, the remaining orthogonal vectors constitute the basis matrix $\hat{\mathbf{U}}$. The projection operator is subsequently formulated as:
$\mathbf{P}=\hat{\mathbf{U}}\hat{\mathbf{U}}^{T}.$ (34)
Through spectral decomposition of $\mathbf{K}_{r}\mathbf{K}_{r}^{T}$, we partition the eigenspace components as follows:
$\mathbf{U}=[\mathbf{U}_{1},\mathbf{U}_{2}],\quad\bm{\Lambda}=\begin{bmatrix}\bm{\Lambda}_{1}&\mathbf{0}\\ \mathbf{0}&\bm{\Lambda}_{2}\end{bmatrix}$, (35)
with $\bm{\Lambda}_{2}$ containing exclusively null eigenvalues and $\mathbf{U}_{2}$ comprising their corresponding eigenvectors. The orthogonality of $\mathbf{U}$ yields:
$\mathbf{U}_{2}^{T}\mathbf{K}_{r}\mathbf{K}_{r}^{T}=\mathbf{U}_{2}^{T}\mathbf{U}_{1}\bm{\Lambda}_{1}\mathbf{U}_{1}^{T}=\mathbf{0}.$ (28)
This establishes that the range space of $\mathbf{U}_{2}$ coincides with the kernel of $\mathbf{K}_{r}\mathbf{K}_{r}^{T}$. Consequently, the projection matrix is equivalently expressed as:
$\mathbf{P}=\mathbf{U}_{2}\mathbf{U}_{2}^{T}.$ (36)
Synthesizing equations (28) and (29), we derive the fundamental property:
$\Delta\mathbf{P}\mathbf{K}_{r}\mathbf{K}_{r}^{T}=\Delta\mathbf{U}_{2}\mathbf{U}_{2}^{T}\mathbf{K}_{r}\mathbf{K}_{r}^{T}=\mathbf{0},$ (37)
confirming that the operator $\Delta\mathbf{P}$ indeed projects any vector $\Delta$ onto the null space of $\mathbf{K}_{r}\mathbf{K}_{r}^{T}$.
C.4 Solution Derivation for Null-space Projection Unlearning
We aim to find the parameter update $\Delta^{*}$ that minimizes the following composite objective function:
$\Delta^{*}=\arg\min_{\Delta}(\underbrace{\left\|(W_{\mathrm{dp}}+\Delta P)K_{f}-V_{f}\right\|^{2}}_{\text{forget term}}+\underbrace{\|\Delta P\|^{2}}_{\text{constraint term}}).$ (38)
First, we set the gradient of the objective function with respect to $\Delta$ to zero:
$\frac{\partial J(\Delta)}{\partial\Delta}=2(W_{\mathrm{dp}}K_{f}-V_{f})K_{f}^{T}P^{T}+2\Delta PK_{f}K_{f}^{T}P^{T}+2\Delta PP^{T}=0.$ (39)
Dividing the entire equation by 2 and rearranging terms to isolate $\Delta$ gives:
$(V_{f}-W_{\mathrm{dp}}K_{f})K_{f}^{T}P^{T}=\Delta P(K_{f}K_{f}^{T}P^{T}+P^{T}).$ (40)
Factoring out $P^{T}$ on the right-hand side results in:
$(V_{f}-W_{\mathrm{dp}}K_{f})K_{f}^{T}P^{T}=\Delta P(K_{f}K_{f}^{T}+I)P^{T}.$ (41)
Assuming $PP^{T}$ is invertible, we can solve for $\Delta$ by right-multiplying both sides by $P(K_{f}K_{f}^{T}+I)^{-1}$, leading to the solution:
$\Delta^{*}=(V_{f}-W_{\mathrm{dp}}K_{f})K_{f}^{T}(K_{f}K_{f}^{T}+I)^{-1}.$ (42)
Finally, by applying the push-through identity $(AB+I)^{-1}A=A(BA+I)^{-1}$ with $A=K_{f}^{T}P$ and $B=K_{f}$, we obtain the elegant and computationally convenient closed-form solution:
$\Delta^{*}=(V_{f}-W_{\mathrm{dp}}K_{f})K_{f}^{T}P(K_{f}K_{f}^{T}P+I)^{-1}.$ (43)
C.5 Multi-layer Udit.